# Fast Hip Joint Moment Estimation with A General Moment Feature Generation Method

### v1

@@ IV-D Generalization of GMF
Table. III presents the performance of GMF in the generalization experiment. Within the scope of this experiment, all four estimators showed an improvement in joint moment estimation accuracy after introducing GMF (compared to the method without body parameters: 7.96% in FFN, 8.12% in CNN, 8.41% in TCN, and compared to the method fusion body parameters but without GMF: 11.44% in FFN, 10.94% in CNN, 9.13% in TCN). Since the GMF generator and decoder used in the experiment were from a pre-trained model and with no parameters change, the results further demonstrate that the generated GMF is a more individual-independent representation of joint moment and contributes to the accuracy of predicting joint moment based on joint kinematics.
V Discussion
@@
This study utilizes a GMF generation method with GRU-based network to achieve a better estimation of hip joint moment. To the best of our knowledge, this is the first study that considers real-time and inter-user variability issues in joint moment estimation and proposes a solution. On one hand, we demonstrate the real-time value of GRU-based neural networks in joint moment estimation. On the other hand, the GMF, which is obtained through an adaptive neural network, leads to significant improvements in model accuracy and has the potential to become an advanced method for future joint moment estimation and analysis.
V-A The Inputs of the Model
@@
Compared to most existing methods for estimating joint moment (using sensor information such as IMU and EMG as input to train the model), the GRU-based estimator used in this paper takes 6-dimensional joint kinematics data as inputs instead of sensor data. This is because the joint moment estimation is considered a model identification problem that requires establishing an approximate model with dynamic characteristics rather than a completely uninterpretable black-box model. For practical applications, model training can be performed using angles obtained from sensors or using highly correlated data of sensors as inputs directly, without requiring changes to the overall model structure. However, additional challenges may be posed in these ways such as sensor accuracy drift and data loss.
@@
Other inputs used by the model are human body parameters, which are composed of height and weight in this paper. These two parameters may not be the optimal inputs for obtaining GMF. Future research could explore more suitable body parameters to achieve better GMF generation and more accurate predictions in the problem of joint moment estimation based on joint kinematics.
V-B Accuracy of the Decoder
To decode the GMF, we use a fully connected neural network of which the structure is the same as the feature generator, and make the loss of decoder back-propagate to the generator with a weighting coefficient to ensure the decodability of the generated GMF. While this method achieves high accuracy (decoding RMSE less than 0.003 Nm/kg) in decoding the GMF into joint moment, the decoding error can only approximate zero rather than reaching it due to the use of a separate neural network. Additionally, to ensure the accuracy of GMF decoding, the decoder may become overly complex, spending extra computation time. This turns what seems like a simple inverse transformation task into a complex and inaccurate one. The concept of reversible neural networks [34, 35, 36], which have lightweight memory consumption and can preserve input information features, has been proposed. When applied to our method, it has the potential to achieve zero error in GMF decoding with fewer parameters. We may further explore the application of such networks in our method in the future.
V-C Time Efficiency Consideration
When comparing the timeliness of estimators, the experimental results shown in this study using a python environment on the same device have clear significance in terms of the relative computation time of the models. However, the absolute values of computation time for each model cannot directly represent the actual computation speed when applied to exoskeletons. To achieve quick response exoskeleton control, methods such as model compression and deployment with C/C++ language are necessary.
@@ V-C Time Efficiency Consideration
Since the proposed method originates from the research on assistive hip exoskeleton control, the estimator model proposed in this paper uses a GRU-based network structure. However, for joint moment estimation tasks that do not require timeliness, the optimal network structure for the estimator may not necessarily be based on recurrent neural networks. The joint moment estimation method we propose can be extended to the field of biomechanics, as demonstrated in the generalization experiment, where it also performs well on mainstream neural networks besides the GRU-based network. For systems that have low real-time requirements, future researchers can use more appropriate network structures within the framework of this method to obtain higher-precision joint moment estimation models.
VI Conclusion
@@
To address the issues of long computation time and lack of personalization in neural networks used for fast joint moment estimation, this study proposes an estimation method based on GMF generation. The method utilizes a generator to adaptively generate GMF, which serves as a more individual-independent representation compared to the biomechanical joint moment. A GRU-based neural network is then employed to predict the generated GMF. The GMF improves the prediction accuracy of the estimator while being decoded into joint moments accurately. Additionally, the GRU-based estimator demonstrates the capability to quickly estimate GMF. Experimental results on prediction accuracy and computation time of the joint moment estimation model validate the effectiveness of the proposed method. In particular, the generalization experiment demonstrates that this method can be applied to various neural networks structure for joint moment estimation, which means it has potential to achieve better accuracy with more refined adjustments to the network structure. The contributions of this work enable better implementation of real-world joint moment estimation and its application in guiding exoskeleton control.

### v2

@@ 4.3 Generalization of GMF
Table. 2 presents the performance of GMF in the generalization experiment. Within the scope of this experiment, all four estimators showed an improvement in joint moment estimation accuracy after introducing GMF (compared to the method without body parameters: 7.96% in FCNN, 8.12% in CNN, 8.41% in TCN, and compared to the method fusion body parameters but without GMF: 11.44% in FCNN, 10.94% in CNN, 9.13% in TCN). Since the GMF generator and decoder used in the experiment were from a pre-trained model and with no parameters change, the results further demonstrate that the generated GMF is a more individual-independent representation of joint moment and contributes to the accuracy of predicting joint moment based on joint kinematics.
Figure 9: Controller performance in variable-speed walking trial. (a) Walking speed varied from 0 m/s to 1.25 m/s, maintained for 10 s, and then decelerated to 0 m/s.(b)-(e) Motor angle, angular velocity, assistance torque, and power trajectories on hip joint with the walking speeds corresponding to (a). The positive direction of the vertical axis represents the flexion of the leg during walking, while the negative direction represents the extension.
Figure 10: (a) Experiment environment of the metabolic cost test. (b) Net metabolic rate for five subjects (*: p$<$0.05, **: p$<$0.01, ***: p$<$0.001). All subjects received the assistance gain compared to Assist Off. (c) Average net metabolic rate for all subjects
4.4 Reducing Metabolism Rate of Users during Walking
Fig. 9 illustrates the assistance performance of the exoskeleton deployed with proposed GMF-based controller, including the assistive torques during three phases: transitioning from standing to walking, steady-state walking at 1.25 m/s, and transitioning from walking to standing. Benefiting from the powerful representational capabilities of neural networks, the GMF model controller can achieve stable assistance during process transitions. The peak torque was about 7.81 Nm for the support phase and 4.29 Nm for the swing phase, respectively. The peak amplitude Difference between the two phases shows the proposed method achieves expected assistance torque as it is widely known the hip joint requires greater moment in the support phase than it in the swing phase during walking. Fig. 9 also shows a more detailed manifestation of the hip joint angle, angular velocity, and power during assistive walking. All of these data were obtained from the motor encoders of hip exoskeleton.
The metabolic cost test results are presented in Fig. 10. During the experiments, the exoskeleton, utilizing the proposed joint moment estimation method, delivers assistive torque to users, effectively reducing their metabolic rate. The mean net metabolic cost reduction achieves 10.5% compared to normal walking (No Exo), while the reduction achieves 20.5% compared to the Assist Off model ($p<0.01$). A notable result is that 2 participants (Subject 3 and Subject 4) in the validation experiments achieved better assistance gain because the participants were most comfortable with the exoskeleton size and had undergone longer walking training for using the exoskeleton, which has been proven to be effective in increasing a subject’s metabolic cost reduction [37].
5 Discussion
@@
This study utilizes a GMF generation method with a GRU-based network to achieve a better estimation of the hip joint moment and its application on the exoskeleton, which benefits from the consideration of inter-user variability in joint moment estimation for exoskeleton assistance. The GMF, which is obtained through an adaptive neural network, leads to significant improvements in model accuracy and has the potential to become an advanced method for future joint moment estimation. The metabolism cost test result showed that our methods achieve the satisfactory results of assistance gain.
5.1 Balance Between Accuracy and Computation Efficiency
This paper proposed a GMF-based neural network framework, aiming to improve estimation accuracy with limited computation sources. To achieve the accuracy superior to the state-of-the-art (SOTA) model, neural networks with large-scale parameters are conventionally required. Our previous research has stated the relationship between accuracy and computation time [27], in which the estimation accuracy of the neural network approaches saturation as the number of network parameters increases. For a real-time exoskeleton control system, efficient computation translates to faster response speed, higher estimation precision, and longer endurance time. This is one of the primary motivations of this study: to explore more efficient estimation methods rather than relying solely on increasing the number of neural network parameters.
It is clear that better estimation accuracy would benefit the exoskeleton control based on instantaneous joint moment estimates. However, there is an uncertain question of whether a better prediction accuracy provides a significant increase in assistance gain. From the results in Table. 2, FCNN, CNN and TCN models have a higher estimation accuracy than the GRU model. However, on the one hand, the other three models have a lower control frequency, which may lead to a poor performance when they were applied on the hip exoskeleton. On the other hand, the multiple comparisons test (with Bonferroni correction) results show that the difference in estimation accuracy among the four models is not completely significant. Therefore, the GRU model is selected as the GMF estimator for the exoskeleton control system in this study.
5.2 Transferability of Using Other Input Sensors
@@
Compared to most of the existing methods for estimating joint moment (using sensor information such as IMU as model input), the GRU-based estimator used in this paper takes joint kinematics data as inputs instead of raw sensor data. Because the joint moment estimation had been considered a model identification problem that requires establishing an approximate model with dynamic characteristics rather than a completely uninterpretable black-box model. This also means that the model controller in this study uses sensor data in a more humanly understandable way, trying to ensure the reliability of the result [38]. Our exoskeleton experiments were able to implement a deep learning-based assistance control model utilizing only motor encoders and got successful results, which we believe it was benefited from this more rational approach to data processing.
Besides, for other potential practical applications, model training can be performed using angles obtained from sensors or using highly correlated data like IMU data of sensors as inputs directly, without requiring changes to the overall model structure. However, additional challenges may be posed in these ways, such as sensor accuracy drift and data transmission loss [39].
5.3 Better Potential Estimation Accuracy
@@
The other important inputs used by the model are human body parameters, which are composed of height and weight in this paper. These two parameters are valid but might not be the optimal attributes for obtaining GMF. Future research could explore more suitable body parameters to achieve better GMF generation and more accurate estimations in the problem of joint moment estimation based on joint kinematics.
@@
Furthermore, since the proposed method originates from the research on assistive hip exoskeleton control, the estimator model proposed in this paper uses a GRU-based network structure. However, for joint moment estimation tasks that do not require timeliness, the optimal network structure for the estimator may not necessarily be based on recurrent neural networks. The joint moment estimation method we propose can be extended to the field of biomechanics, as demonstrated in the generalization experiment, where it also performs well on mainstream neural networks besides the GRU-based network. For systems with low real-time requirements, future researchers can use more appropriate network structures within the framework of this method to obtain higher-precision joint moment estimation models.
5.4 Limitations
The proposed method still has several limitations that need to be further addressed in future research.
5.4.1 Burden of Wearing Exoskeleton
Fig. 10 has shown a 12.5% metabolic cost increase when subjects wearing the hip exoskeleton without power compared to normal walking. Although this study focused on achieving accurate joint moment estimation, there is a clear need for exoskeleton structures with low wearing burden to enhance the assistance gain for users. Meanwhile, as the results of the experiment more conclusively verified that training benefits the exoskeleton user to reach a higher assistance gain, future studies may need to consider how to control this impact in experiments.
5.4.2 Accuracy of the Decoder
To decode the GMF, we use a fully connected neural network with the same structure as the feature generator, and make the loss of decoder back-propagate to the generator with a weighting coefficient to ensure the decodability of the generated GMF. While this method achieves high accuracy (decoding RMSE less than 0.003 Nm/kg) in decoding the GMF into the joint moment, the decoding error can only approximate zero rather than reaching it due to the use of a separate neural network. Additionally, to ensure accurate GMF decoding, the decoder may become overly complex, leading to increased computational time. This complexity transforms what should be a straightforward inverse transformation into a more intricate and less reliable process. The concept of reversible neural networks [40, 41, 42], which have lightweight memory consumption and can preserve input information features, has been proposed. When applied to our method, it has the potential to achieve zero error in GMF decoding with fewer parameters. We may further explore the application of such networks in our method in the future.
5.4.3 Challanges of Exoskeleton Control with Encoder-only Sensors
Inspired by the idea that any sensor capable of measuring joint kinematics can be used to estimate joint moments, we implemented exoskeleton control using motor encoders alone. However, encoder-only sensors still face significant limitations in exoskeleton applications, primarily due to the known issue of coupling between motor motion and human movement.
Our experiments show that encoder-only sensors can provide stable and accurate estimations of joint moments when the assistance time is appropriately calibrated and comfortable for the user. However, determining the correct assistance timing requires pre-experimentation to identify the optimal delay time factor. Furthermore, if the assistance time is not accurately adjusted, the assistive ratio can significantly amplify the negative effects, leading to suboptimal performance. To address this issue in the future, more specific methods should be developed, such as the use of Series Elastic Actuators (SEA) [9], abnormal assistance detection techniques, and other potential solutions.
6 Conclusion
@@
To address the issues of lack of personalization joint moment estimation neural network for hip exoskeleton assistance, this study proposes an estimation method based on GMF generation. The method utilizes a generator to adaptively generate GMF, which serves as a more individual-independent representation compared to the biomechanical joint moment. A GRU-based neural network is then employed to estimate the generated GMF. The GMF improves the estimation accuracy of the estimator while could be decoded into joint moments accurately. Experimental results demonstrate enhanced estimation accuracy and strong generalization capability of the proposed method. Furthermore, real-world exoskeleton tests using encoder-only sensors validate its effectiveness in delivering assistive benefits to users. The contributions of this work advance the feasibility of real-time joint moment estimation and its practical application in exoskeleton control.
Acknowledgments
We sincerely thank all the students in HARLab for their invaluable support of this work, as well as all the participants in our experiments for their time and effort.
Author Contributions
Yuquan Leng and Chenlong Fu supervised the project. Yuanwen Zhang and Jingfeng Xiong conceived the idea and designed the experiments. Chuheng Chen and Xinxing Chen reviewed the manuscript and proposed refinements of the experiment. All authors reviewed the manuscript and approved the version to be submitted.
Funding
This work was supported by National Key R&D Program of China [grant numbers 2024YFC3082800]; the National Natural Science Foundation of China [grant numbers 52175272]; Guangdong Basic and Applied Basic Research Foundation [grant numbers 2024B1515020008 and 2023B1515130007]; and Shenzhen Science and Technology Program [grant numbers KCXFZ20230731093401004, RCYX20231211090345058 and JCYJ20220530114809021].
Declaration of interest
None
Data Availability
All data needed to evaluate the conclusions are presented in the manuscript and/or the Supplementary Materials. Additional data that support the findings of this study are available from the corresponding author upon reasonable request
Supplementary Materials
Movies S1 and S2.